# R-LLaVA: Improving Med-VQA Understanding through Visual Region of Interest

### v2

@@ Abstract
R-LLaVA: Improving Med-VQA Understanding
through Visual Region of Interest
@@
1 Introduction
Medical Visual Question Answering (Med-VQA) has recently garnered significant attention Chen et al. (2022b); Gong et al. (2021); Ren and Zhou (2020); Khare et al. (2021). As an emerging area in medical AI, Med-VQA aims to answer medical questions in natural language based on input medical images. A robust Med-VQA system can assist clinicians in interpreting medical images, thereby ensuring accuracy and expediting the diagnostic process. For patients, automated Med-VQA services can greatly meet the demand for personalized healthcare consultations Liu et al. (2023a).
Numerous deep learning-based approaches have been explored in the realm of Med-VQA Tiong et al. (2022); Banerjee et al. (2021); Changpinyo et al. (2022); Liu et al. (2023b); Gai et al. (2024). For instance, Nguyen et al. (2019) utilized Bilinear Attention Networks (BAN) Kim et al. (2018), enhancing them with Mixed Enhanced Visual Features (MEVF), which integrates pre-trained meta-learning modules and Convolutional Denoising Autoencoders (CDAE) to improve the performance of Med-VQA models. Building on this, Zhan et al. (2020) proposed a conditional reasoning framework to further enhance the inference capabilities of Med-VQA models. However, many of these methods tend to underperform in practical scenarios, primarily due to limitations in extracting and integrating information from a limited number of medical images and text data Eslami et al. (2021); Song et al. (2022); Wang et al. (2022). To address this, Eslami et al. (2021) introduced the CLIP architecture into the MEVF framework Nguyen et al. (2019), using CLIP as the visual encoder pre-trained on the multimodal medical dataset ROCO Pelka et al. (2018), which demonstrated significant performance gains. Additionally, Liu et al. (2023a) developed VQA-Adapter, a lightweight adapter that, combined with label smoothing, efficiently fine-tunes CLIP for Med-VQA, reducing computational costs and mitigating overfitting. Li et al. (2024) proposed LLaVA-Med, leveraging GPT-4 and a novel curriculum learning approach to efficiently train LLaVA on biomedical images, further enhancing Med-VQA’s capabilities.
However, previous Med-VQA approaches have often overlooked the importance of medical and spatial priors that are essential for disease localization in Med-VQA tasks. Inspired by clinical practices where doctors rely on such priors, we propose the R-LLaVA framework, which incorporates doctors’ domain knowledge and spatial annotations into Med-VQA models. Fig. 1 shows the major idea of the proposed method. Specifically, we integrate simple physician annotations, such as bounding boxes (bbox), as prior knowledge, directly injecting these visual cues into the image space via CLIP showing CLIP’s ability to interpret visual markers). These annotated regions are then fed into the LLaVA model during training. Our experiments show that even minimal annotations from doctors can significantly improve the accuracy of the model. R-LLaVA consistently outperforms state-of-the-art Med-VQA methods across four large-scale datasets, demonstrating remarkable performance improvements. Our work has two major contributions:
  • The work introduces R-LLaVA method to improve Med-VQA by incorporating regions of visual interest through a three-phase training process.
@@ 2.1 Medical Dataset Reconstruction with Region-of-Interest (RoI) VQA
As is shown in Fig. 2, the reconstructed dataset comprises four types of QA pairs: (1) Region localization - where the model is required to predict the bounding box coordinates corresponding to a described region, e.g., ”Please provide the bounding box coordinate of the region this sentence describes: Heart". (2) Region selection - where the model is required to select the bounding box among four corresponding to a described region, e.g., "Select the bounding box (bbox) describes spleen. A. Yellow B. Purple C. Green D. Red". (3) Region description with bounding box coordinates - where the model is asked to provide a description given a bounding box, e.g., "Please provide a short description for this region: [115, 163, 243, 268]". (4) Region description with Bounding Box highlighted in input image - where the model is asked to describe the bounding box, e.g., "Please provide a short description inside the bounding box".
@@
2.2 R-LLaVA Training Stage I: Pretraining
The first stage of training is to align the biomedical concepts while maintaining efficiency. As shown in Fig. 3(a), the image encoder (e.g., CLIP (Radford et al., 2021)) and the language model (e.g., Vicuna (Zheng et al., 2023)) are frozen in stage I. Only the multimodal connector is trained by updating the projection matrix. This freezing technique is crucial for efficiency and understanding concepts, as it allows the model to connect the foundational representations of visual and textual embeddings under medical settings. The major dataset used at this stage follows the LLaVA-Med alignment dataset Li et al. (2023). It is built upon a large-scale dataset of 15 million parallel figure-caption pairs for biomedical vision-language processing—PMC-15M. PMC-15M is shrunk to only 600,000 pairs to balance concept coverage and efficiency, which speeds up training while still providing great alignment performance.
Figure 3: R-LLaVA Training Pipeline (Stage I and Stage II)
@@ 2.3 R-LLaVA Training Stage II: Instruction Tuning with Visual RoI Approach
where $\alpha$ controls the transparency level, $\mathbf{X}_{\mathrm{img}}$ is the base image, and $\mathbf{P}_{\mathrm{doc}}$ represents the medical regions of interest. The composite image $\bar{\mathbf{X}}_{\mathrm{merged}}$ is then fed into the multimodal language model to guide its attention toward these key areas.
To capture both detailed and abstract information, we extract multi-level features from several layers of CLIP. The shallower layer is used for capturing finer details, while deeper layers capture more abstract semantic representations (Ruan et al., 2024). These features are combined, normalized using LayerNorm for stability, and then processed by an MLP layer to integrate the diverse visual cues effectively.
@@
$\begin{split}&P(\mathbf{Y}\mid\bar{\mathbf{X}}_{\mathrm{merged}},\mathbf{X}_{\text{instruct}})\\ &=\prod_{t=1}^{T}P_{\boldsymbol{\Omega}}(x_{i}\mid\bar{\mathbf{X}}_{\mathrm{merged}},\mathbf{X}_{\text{instruct}},\mathbf{Y}_{<t})\end{split}$ (1)
where $\bar{\mathbf{X}}_{\mathrm{merged}}$ contains the integrated images with an emphasis on regions of interest, $\mathbf{X}_{\text{instruct}}$ provides the textual instruction, which can be constructed from questions. $T$ is the sequence length of the target answer $\mathbf{Y}$, $\mathbf{Y}_{<t}$ includes all the answer tokens preceding the current prediction token $x_{t}$, and $\boldsymbol{\Omega}$ represents the trainable parameters.
@@ 3 Experiment
In this section, we first provide a detailed description of the utilized dataset (Section 3.1), including the types of questions it encompasses. Next, we present the selected evaluation metrics (Section 3.2). Based on the chosen datasets and metrics, we outline the training strategies and settings employed in our experiments (Section 3.3). We present the main results in Section 3.4, along with a comprehensive ablation study analyzing the impact of our methodological choices (Section 3.5). Additional qualitative examples are provided in the Appendix for further reference.
@@ 3.1 Dataset
For pretraining, we utilize a large-scale medical image-caption dataset from Li et al. (2024), containing 600K pairs to facilitate foundational alignment with medical concepts. For Fine-tuning, we employ four VQA datasets: VQA-RAD, SLAKE-EN, PathVQA, and VQA-Med. We augment the SLAKE-EN dataset with RoI QA, as detailed in Section 2.1.
@@ 3.3 Training Strategy and Setting
Both pretraining and fine-tuning stages are conducted on 8 A100 80G GPUs. Pretraining requires approximately 6 hours for the 7B model and 12 hours for the 13B model, while fine-tuning takes 3 hours for the 7B model and 6 hours for the 13B model.
@@ 3.5.3 Visual Region-of-Interest (RoI) Approach
Alpha Blending Strategy This experiment analyzes the impact of varying bounding box opacity (alpha) on VQA performance across multiple datasets. Higher alpha values represent a more visible bounding box, while lower values signify a subtler presence. Alpha values include fixed levels (96, 128, 255) and a randomly sampled range within [96, 255]. For all VQA datasets, alpha values of 96, 128, and 255 yield similar results, with slight incremental improvements as alpha increases. This suggests that a more visible bounding box can aid in certain tasks but may not significantly influence outcomes. Randomized Alpha within the range [96, 255] achieves the best results across all VQA datasets compared to other fixed alpha groups. The variability introduced by random alpha values likely provides a more robust learning experience by exposing the model to diverse bounding box appearances, resulting in improved generalization and accuracy.
@@ 4.1 Visual Large Language Model
Integrating visual understanding into large language models has significantly advanced multimodal AI capabilities. Early models like ViLBERT Lu et al. (2019) and VisualBERT Li et al. (2019) extended the BERT architecture to process both text and images, enabling tasks such as visual question answering and image captioning. Contrastive learning frameworks like CLIP Radford et al. (2021) learned joint representations of images and text from large-scale datasets, achieving impressive zero-shot performance on various tasks. Recent developments focus on a more seamless integration of vision and language. Models like Flamingo Alayrac et al. (2022) and PaLI Chen et al. (2022a) incorporate visual information into language models using gated cross-attention and scale-up pre-training with multilingual data. Furthermore, the enhanced multimodal features of ChatGPT and Gemini Team et al. (2023) model represent significant strides toward more integrated and capable AI systems.
@@ 4.2 Med-VQA
Med-VQA is a method where a model answers questions from patients or clinicians based on medical images like CT scans, MRI scans, or pathology images. With the advancement of deep learning, researchers have proposed numerous Med-VQA methods. M2I2  Li et al. (2022b) is a self-supervised vision-language pretraining method that significantly improves medical VQA performance by learning multimodal representations through masked modeling and contrastive learning. However, its ability to predict free-form answers may be influenced by the complexity of the dataset. BiomedCLIP Zhang et al. (2024) is a multimodal biomedical foundation model trained on PMC-15M. It outperforms earlier vision-language models like PubMedCLIP Eslami et al. (2023b) and MedCLIP, as well as radiology-specific models such as BioViL Boecking et al. (2022). Bazi  Bazi et al. (2023a) proposes a vision-language model based on a Transformer encoder-decoder architecture. It leverages the CLIP model for semantic embedding and uses a generative decoder to produce answers in an autoregressive manner. Recently, inspired by LLaVA (Liu et al., 2024), LLaVA-Med (Li et al., 2024) fine-tunes LLaVA by self-generated biomedical instruction-following dataset to address challenges in medical image interpretation. However, the explicit handling of region-specific information in complex medical scenarios remains insufficiently explored.
@@ 5 Conclusion
The proposed method, which builds on LLaVA (Liu et al., 2024; Cai et al., 2024), inherits several inherent challenges associated with VLLMs, such as hallucination and biases. Regarding hallucination, as demonstrated in Table 1 and Fig. 4, R-LLaVA makes more effective use of the visual regions of interest, leading to generally superior performance. Nevertheless, it may still produce responses that are not grounded in factual information or the input data. We consider this work a significant step toward enhancing biomedical VQA by utilizing doctor guidance. To address potential biases (Ruan et al., 2023), particularly those that could favor or disfavor specific demographics, we test R-LLaVA on cases where annotations were unavailable due to privacy constraints. Additionally, integrating improved vision or language encoders could further alleviate these biases. Given the high computational cost and energy demand of training LLMs in general, we leverage pre-trained image and language encoders. This approach is especially appropriate given the relatively small size of biomedical datasets, eliminating the need for training from scratch.
@@
Appendix A Qualitative Examples
Figure 5: More qualitative comparison of medical visual question answering from SLAKE dataset.
Figure 6: More qualitative comparison of medical visual question answering from SLAKE dataset.
Figure 7: More qualitative comparison of medical visual question answering from SLAKE dataset.

### v5

@@ Abstract
1 Introduction
Medical Visual Question Answering (Med-VQA) has recently garnered significant attention Chen et al. (2022b); Gong et al. (2021); Ren & Zhou (2020); Khare et al. (2021). As an emerging area in medical AI, Med-VQA aims to answer medical questions in natural language based on input medical images. A robust Med-VQA system can assist clinicians in interpreting medical images, thereby ensuring accuracy and expediting the diagnostic process. For patients, automated Med-VQA services can greatly meet the demand for personalized healthcare consultations Liu et al. (2023a).
Numerous deep learning-based approaches have been explored in the realm of Med-VQA Tiong et al. (2022); Banerjee et al. (2021); Changpinyo et al. (2022); Liu et al. (2023b); Gai et al. (2024); Liu et al. (2024c). For instance, Nguyen et al. (2019) utilized Bilinear Attention Networks (BAN) Kim et al. (2018), enhancing them with Mixed Enhanced Visual Features (MEVF), which integrates pre-trained meta-learning modules and Convolutional Denoising Autoencoders (CDAE) to improve the performance of Med-VQA models. Building on this, Zhan et al. (2020) proposed a conditional reasoning framework to further enhance the inference capabilities of Med-VQA models. However, many of these methods tend to perform poorly in practical scenarios, mainly due to limitations in the extraction and integration of information from a limited number of medical images and text data Eslami et al. (2021); Song et al. (2022); Wang et al. (2022); Liu et al. (2024b; 2025). To address this, Eslami et al. (2021) introduced the CLIP architecture into the MEVF framework Nguyen et al. (2019), using CLIP as the visual encoder pre-trained in the ROCO multimodal medical dataset Pelka et al. (2018), which demonstrated significant performance gains. Additionally, Liu et al. (2023a) developed VQA-Adapter, a lightweight adapter that, combined with label smoothing, efficiently fine-tunes CLIP for Med-VQA, reducing computational costs and mitigating overfitting. Li et al. (2024) proposed LLaVA-Med, leveraging GPT-4 and a novel curriculum learning approach to efficiently train LLaVA in biomedical images, further enhancing the capabilities of Med-VQA.
@@ 1 Introduction
However, previous Med-VQA approaches have often overlooked the importance of medical and spatial priors that are essential for disease localization in Med-VQA tasks. Inspired by clinical practices in which doctors rely on such priors, we propose the R-LLaVA framework, which incorporates doctors’ domain knowledge and spatial annotations into Med-VQA models. Fig. 1 shows the main idea of the proposed method. Specifically, we integrate simple physician annotations, such as bounding boxes (bbox), as prior knowledge, directly injecting these visual cues into the image space via CLIP showing CLIP’s ability to interpret visual markers. These annotated regions are then fed into the LLaVA model during training. Our experiments show that even minimal annotations from doctors can significantly improve the accuracy of the model. R-LLaVA consistently outperforms the state-of-the-art Med-VQA methods in four large-scale datasets, demonstrating remarkable performance improvements. Our work has two major contributions:
  • The work introduces R-LLaVA method to improve Med-VQA by incorporating regions of visual interest through a two-phase training process.
@@ 2.1 Medical Dataset Reconstruction with Region-of-Interest (RoI) VQA
As is shown in Fig. 2, the reconstructed dataset comprises four types of QA pairs: (1) Region localization, where the model is required to predict the bounding box coordinates corresponding to a described region, e.g., "Please provide the bounding box coordinate of the region this sentence describes: Heart". (2) Region selection - where the model is required to select the bounding box among four corresponding to a described region, e.g., "Select the bounding box (bbox) describes spleen. A. Yellow B. Purple C. Green D. Red". (3) Region description with bounding box coordinates - where the model is asked to provide a description given a bounding box, e.g., "Please provide a short description for this region: [115, 163, 243, 268]". (4) Region description with Bounding Box highlighted in input image, where the model is asked to describe the bounding box, e.g., "Please provide a short description inside the bounding box".
@@
2.2 R-LLaVA Training Stage I: Pre-training
The first stage of training is to align the biomedical concepts while maintaining efficiency. As shown in Fig. 3(a), the image encoder (e.g., CLIP (Radford et al., 2021)) and the language model (e.g. Vicuna (Zheng et al., 2023)) are frozen in stage I. Only the multi-modal connector is trained by updating the projection matrix. This freezing technique is crucial for efficiency and understanding concepts, as it allows the model to connect the foundational representations of visual and textual embeddings under medical settings. The major dataset used at this stage follows the LLaVA-Med alignment dataset Li et al. (2023). It is built upon a large-scale dataset of 15 million parallel figure-caption pairs for biomedical vision-language processing—PMC-15M. PMC-15M is shrunk to only 600,000 pairs to balance concept coverage and efficiency, which speeds up training while still providing great alignment performance.
@@ 2.3 R-LLaVA Training Stage II: Instruction Tuning with Visual RoI Approach
where $\alpha$ controls the transparency level, $\mathbf{X}_{\mathrm{img}}$ is the base image, and $\mathbf{P}_{\mathrm{doc}}$ represents the medical regions of interest. The composite image $\bar{\mathbf{X}}_{\mathrm{merged}}$ is then fed into the multimodal language model to guide its attention toward these key areas Vaswani et al. (2017); Dosovitskiy et al. (2021); Ruan & Di (2024).
To capture both detailed and abstract information, we extract multi-level features from several layers of CLIP. The shallower layer is used for capturing finer details, while deeper layers capture more abstract semantic representations (Ruan et al., 2024a). These features are combined, normalized using Layer Normalization for stability, and then processed by a Multi-Layer Perceptron(MLP) layer to integrate the diverse visual cues effectively.
Figure 3: R-LLaVA Training Pipeline (Stage I and Stage II)
@@
$\begin{split}&P(\mathbf{Y}\mid\bar{\mathbf{X}}_{\mathrm{merged}},\mathbf{X}_{\text{instruct}})\\ &=\prod_{t=1}^{T}P_{\bm{\Omega}}(x_{t}\mid\bar{\mathbf{X}}_{\mathrm{merged}},\mathbf{X}_{\text{instruct}},\mathbf{Y}_{<t})\end{split}$ (1)
where $\bar{\mathbf{X}}_{\mathrm{merged}}$ contains the integrated images with an emphasis on regions of interest, $\mathbf{X}_{\text{instruct}}$ provides the textual instruction, which can be constructed from questions. $T$ is the sequence length of the target answer $\mathbf{Y}$, $\mathbf{Y}_{<t}$ includes all the answer tokens preceding the current prediction token $x_{t}$, and $\bm{\Omega}$ represents the trainable parameters.
@@ 3 Experiment
In this section, we first provide a detailed description of the utilized dataset (Section 3.1), including the types of questions it encompasses. Next, we present the selected evaluation metrics (Section 3.2). Based on the chosen datasets and metrics, we outline the training strategies and settings employed in our experiments (Section 3.3). We present the main results in Section 3.4, along with a comprehensive ablation study analyzing the impact of our methodological choices (Section 3.5).
@@ 3.1 Dataset
For pretraining, we utilize a large-scale medical image-caption dataset from Li et al. (2024), containing 600K pairs to facilitate foundational alignment with medical concepts. For Fine-tuning, we employ four VQA datasets: VQA-RAD, SLAKE-EN, PathVQA, and VQA-Med. We augment the SLAKE-EN dataset with RoI QA, as detailed in Section 2.1. We adopt a consistent 80/20 train-test split for all datasets, including VQA-RAD, SLAKE-EN, PathVQA, and VQA-Med 2019.
@@ 3.3 Training Strategy and Setting
Both pretraining and fine-tuning stages are conducted on 8 A100 80G GPUs. Pretraining requires approximately 6 hours for the 7B model and 12 hours for the 13B model, while fine-tuning takes 3 hours for the 7B model and 6 hours for the 13B model. All models use FP16 precision for both training and inference.
@@ 3.5.3 Visual Region-of-Interest (RoI) Approach
Alpha Blending Strategy This experiment analyzes the impact of varying bounding box opacity (alpha) on VQA performance across multiple datasets. Higher alpha values represent a more visible bounding box, while lower values signify a subtler presence. Alpha values include fixed levels (96, 128, 255) and a randomly sampled range within [96, 255]. For all VQA datasets, alpha values of 96, 128, and 255 yield similar results, with slight incremental improvements as alpha increases. This suggests that a more visible bounding box can aid in certain tasks but may not significantly influence outcomes. Randomized Alpha (Dynamic) within the range [96, 255] achieves the best results across all VQA datasets compared to other fixed alpha groups. The variability introduced by random alpha values likely provides a more robust learning experience by exposing the model to diverse bounding box appearances, resulting in improved generalization and accuracy.
@@ 4.1 Visual Large Language Model
Recent years have witnessed significant advancements in Large Language Models (LLMs) Brown et al. (2020). Concurrently, developments in Large Vision Models (LVMs) have emerged, complementing LLMs through their capacity for visual understanding. Integrating multimodal understanding into large language models has significantly advanced multimodal AI capabilities Ruan et al. (2024b; 2025). Early models like ViLBERT Lu et al. (2019) and VisualBERT Li et al. (2019) extended the BERT architecture to process both text and images, enabling tasks such as visual question answering and image captioning. Contrastive learning frameworks like CLIP Radford et al. (2021) learned joint representations of images and text from large-scale datasets, achieving impressive zero-shot performance on various tasks. Recent developments focus on a more seamless integration of vision and language. Models like Flamingo Alayrac et al. (2022) and PaLI Chen et al. (2022a) incorporate visual information into language models using gated cross-attention and scale-up pre-training with multilingual data. Furthermore, the enhanced multimodal features of ChatGPT and Gemini Team et al. (2023) model represent significant strides toward more integrated and capable AI systems.
@@ 4.2 Med-VQA
Med-VQA is a method where a model answers questions from patients or clinicians based on medical images like CT scans, MRI scans, or pathology images. With the advancement of deep learning, researchers have proposed numerous Med-VQA methods. M2I2  Li et al. (2022b) is a self-supervised vision-language pretraining method that significantly improves medical VQA performance by learning multimodal representations through masked modeling and contrastive learning. However, its ability to predict free-form answers may be influenced by the complexity of the dataset. BiomedCLIP Zhang et al. (2024) is a multimodal biomedical foundation model trained on PMC-15M. It outperforms earlier vision-language models like PubMedCLIP Eslami et al. (2023a) and MedCLIP, as well as radiology-specific models such as BioViL Boecking et al. (2022). Bazi  Bazi et al. (2023a) proposes a vision-language model based on a Transformer encoder-decoder architecture. It leverages the CLIP model for semantic embedding and uses a generative decoder to produce answers in an autoregressive manner. Recently, inspired by LLaVA (Liu et al., 2024a), LLaVA-Med (Li et al., 2024) fine-tunes LLaVA by self-generated biomedical instruction-following dataset to address challenges in medical image interpretation. However, the explicit handling of region-specific information in complex medical scenarios remains insufficiently explored.
@@ 5 Conclusion
The proposed method, which builds on LLaVA (Liu et al., 2024a; Cai et al., 2024), inherits several inherent challenges associated with VLLMs, such as hallucination and biases. Regarding hallucination, as demonstrated in Table 1 and Fig. 4, R-LLaVA makes more effective use of the visual regions of interest, leading to generally superior performance. Nevertheless, it may still produce responses that are not grounded in factual information or the input data. We consider this work a significant step toward enhancing biomedical VQA by utilizing doctor guidance. To address potential biases (Ruan et al., 2023), particularly those that could favor or disfavor specific demographics, we test R-LLaVA on cases where annotations were unavailable due to privacy constraints. Additionally, integrating improved vision or language encoders could further alleviate these biases. Given the high computational cost and energy demand of training LLMs in general, we leverage pre-trained image and language encoders. This approach is especially appropriate given the relatively small size of biomedical datasets, eliminating the need for training from scratch.